# The “Little Dark Dot”: Evidence for Self-Interacting Dark Matter in the Strong Lens SDSSJ0946+1006?

### v1

@@ 2.2 Computing the full time evolution of the SIDM halo
Figure 2: The density and mass profiles of a core-collapsed SIDM halo with $M_{200}=10^{11}~{}{\rm M_{\odot}}$ (corresponding to the evolutionary stage indicated by the red star in Fig. 1) and its CDM counterpart. SIDM and CDM profiles are shown in blue and black, respectively. Left: 3D density profiles, with the blue vertical dash-dotted line marking the stitching radius, $r_{1}$. Middle: 2D projected surface density profiles. Right: 2D enclosed mass profiles.
@@
In an SIDM halo, heat transport driven by collisions between dark matter particles governs the temporal evolution of halo internal structures. Initially, these collisions facilitate heat flow inward along a positive temperature gradient, leading to the formation of a central core. This process, known as core-formation, results in the gradual growth of the core as the inner region undergoes thermalization. Once thermal contact extends to the outskirts of the halo, heat begins to flow outward, marking the onset of the core-collapse phase. Due to the negative specific heat characteristic, the core area becomes hotter and denser after losing heat, leading the temperature and density of the isothermal core at the halo center increase steadily, culminating in a gravothermal catastrophe (S. Tulin & H.-B. Yu, 2018).
@@
The full time evolution of an SIDM halo can be more rigorously captured using the gravothermal fluid formalism, which models SIDM halos by solving a set of coupled partial differential equations for matter and heat transport with a CDM halo as the initial condition (e.g. R. Essig et al., 2019; H. Nishikawa et al., 2020; S. Yang et al., 2023). In contrast, the isothermal Jeans method assumes that the inner region is in approximate equilibrium, an assumption that breaks down during the core-collapse phase, resulting the standard approach becomes inadequate for capturing the behavior of SIDM halos in this regime.
@@
However, F. Jiang et al. (2023) noted that prior to core-collapse, the isothermal Jeans method allows for both high-density and low-density solutions for the isothermal core: during the minimization of $\delta^{2}$, two sets of solutions $(\rho_{0},\nu_{0})$ are found, where the $\nu_{0}$ values are close to each other, but the $\rho_{0}$ values differ by several orders of magnitude (refers to Figure 1 of F. Jiang et al. 2023). These two sets of solutions gradually converge at a specific time, denoted as $t_{\rm merge}$.
@@
Based on this, S. Yang et al. (2024, inspired by private communications with M. Kaplinghat) demonstrated that mirroring the high-density solution at the merger time $t_{\rm merge}$ yields isothermal solutions corresponding to the core-collapse phase. This leads to a natural definition of the collapse timescale as $t_{\rm coll}=2\times t_{\rm merge}$. As shown in Figure 2 of S. Yang et al. (2024), although minor discrepancies exist, the gravothermal and isothermal solutions exhibit remarkable similarity throughout the entire halo evolution. The physical justification for this mirroring approach can be understood as follows: Since time is treated as a scalar quantity in Eq. 5, the formulation inherently lacks temporal directionality. This means the isothermal Jeans model invariably solves for density profiles that best approximate an isothermal sphere configuration at any given instant – even during-core collapse. By mirroring the high-density solution about $t_{\rm merge}$, we essentially reconstruct the temporal sequence of this evolutionary phase while recovering the proper chronological ordering.
@@
In this work, we adopt the “mirror” strategy to calculate density profile of SIDM halo in core-collapse state. For a given CDM halo with known $M_{200}$ and $c_{200}$, we first compute the time $t_{\rm merge}\cdot\sigma_{m}$, in units of ${\rm Gyr\cdot cm^{2}/g}$, when the high-density and low-density solutions merge. Then, by progressively increasing $t_{\rm age}\cdot\sigma_{m}$ values in Eq. 5 and solving for the density profile of the isothermal core, we determine which solution to adopt: if $t_{\rm age}\cdot\sigma_{m}<t_{\rm merge}\cdot\sigma_{m}$, the low-density solution is used; otherwise, the mirrored high-density solution is applied. This process is illustrated in Fig. 1 for a halo with $M_{200}=1\times 10^{11}~{}{\rm M_{\odot}}$, where the $c_{200}$ is calculated using the mass-concentration relation for CDM from A. A. Dutton & A. V. Macciò (2014).
@@
In the right subplot of Fig. 1, we present the evolutionary scenario of the velocity dispersion, $\nu_{0}$, combining the low-density solution and the mirrored high-density solution, omitting the original high-density solution for clarity. After its initial increase, $\nu_{0}$ remains nearly constant for a long time before $t_{\rm merge}$ but increases steadily afterward. This behavior underscores the rationale for using $t_{\rm merge}$ as the starting point for the core-collapse phase in the isothermal Jeans method.
@@
In Fig. 2 we compare the density and mass profiles of a core-collapsed SIDM halo ($M_{200}=10^{11}~{}{\rm M_{\odot}}$) with its CDM counterpart, highlighting differences in matter distribution. The corresponding evolutionary stage is marked by a red star in Fig. 1. The inner region of the SIDM halo develops a steep and compact core with a radius smaller than 1 kpc, while the density profiles of the SIDM and CDM halos converge at $r_{1}$. The core-collapse-induced inward mass aggregation results in the SIDM halo exhibiting higher densities than the CDM halo within the 0.03 to 3 kpc range. This steeper inner density profile enables the SIDM halo to accumulate more projected mass up to several kpc, with the increase in enclosed projected mass being most pronounced at 1 kpc. It also leads to a steeper slope of the surface density around 1 kpc, as illustrated in the middle and right panel of Fig. 2.
@@ 2.3 Computing the Lensing Properties
To compare with observations of J0946+1006, we calculate two lensing properties which, following Q. Minor et al. (2021), have been identified as well-constrained through lens modeling: (1) the average logarithmic slope of the surface density profile within 1 kpc, denoted as $\gamma_{\rm 2D}~{}(1~{}{\rm kpc})$, and (2) the projected total mass within 1 kpc, $M_{\rm 2D,tot}~{}(<1~{}{\rm kpc})$.
We adopt these two properties from the best-fit results of “tNFWmult” model in Q. Minor et al. (2021), along with model results from G. Despali et al. (2024), as observational benchmarks for comparison with our SIDM model results. To assess whether the modeled SIDM halos produce observational signatures akin to those detected, and to determine the evolutionary stages at which these occur, we computed their complete time evolution and determined the density profiles at specific snapshots along their evolutionary paths, as illustrated by the black dots in Fig. 1.
@@
To measure $\gamma_{\rm 2D}$ at 1 kpc, we fit the surface mass density profile $\Sigma(R)$ with a power-law function $\Sigma(R)=\Sigma_{0}R^{\gamma}$ over the radial range 0.75–1.25 kpc, matching the analysis region of Q. Minor et al. (2021). $M_{\rm 2D}~{}(<1~{}{\rm kpc})$ is then obtained by integrating $\Sigma(R)$ from 0.001 to 1 kpc along $R$.
@@ 3.1 Comparison with Observations
In Fig. 3, we compare the lensing properties of SIDM halos, evaluated after $t_{\rm merge}$, with those of the observed dark perturber in J0946+1006 (Q. Minor et al., 2021; G. Despali et al., 2024). We consider four halo masses in ascending order: $5\times 10^{9}$, $3\times 10^{10}$, $1\times 10^{11}$, and $5\times 10^{11}~{}M_{\odot}$. Their $c_{200}$ values are calculated using the mass-concentration relation from A. A. Dutton & A. V. Macciò (2014).
Our analysis indicates that, across all four selected halo masses, the matter condensation resulting from core-collapse causes $\gamma_{\rm 2D}~{}(1~{}{\rm kpc})$ to reach approximately -1.4 at its steepest point, which is steeper than the singular isothermal sphere case, -1. This steepening brings the slope close to the value reported by Q. Minor et al. (2021): $\gamma_{\rm 2D}~{}(1~{}{\rm kpc})=-1.27^{+0.11}_{-0.13}$. However, only the core-collapsed halo with a mass of $10^{11}~{}{\rm M_{\odot}}$ achieves an enclosed mass comparable to the observed value of $(3.3\pm 0.3)\times 10^{9}~{}{\rm M_{\odot}}$. Halos with masses significantly larger or smaller than the optimal range either accumulate too much or too little mass in this phase, inconsistent with the observation constraints.
@@
Another important observation is that, for a given halo mass, $M_{\rm 2D}~{}{\rm(<1~{}kpc)}$ gradually converges to a finite value during the core-collapse stage, and after $\gamma_{\rm 2D}~{}{\rm(1~{}kpc)}$ reaches its turning point, it reverses accordingly. This is because the dark matter self-interactions is only effective at the inner region of the halo, and the redistributable mass during core collapse is constrained by the total mass of the dark halo within $r_{1}$. Therefore, for any given CDM halo mass, the properties of the density core created by SIDM are not arbitrary; they depend on the initial density profile of the halo at the beginning of its evolution.
@@
From another perspective, the degree of core-collapse in the SIDM halo can be analyzed within the gravothermal fluid framework. In this context, radial heat transport, driven by dark matter self-interactions, can be characterized by the ratio of the scattering mean free path to the halo’s dynamical scale height. The time at which the innermost shell of an SIDM halo transitions to the short mean-free-path regime signifies the onset of post-self-similar collapse. At this point, the inner density begins to increase dramatically, making this time a defining collapse timescale, denoted as $t_{\rm c}$. R. Essig et al. (2019) performed numerical simulations and obtained numerical templates between the core-collapse timescale and other model parameters:
@@
where $\beta=0.75$ is calibrated with $N$-body simulations. Based on this scaling relation, we can transform the product $t_{\rm age}\cdot\sigma_{m}$ into a dimensionless time variable $\tau=t/t_{\rm c}$. This allows us to analyze the degree of core-collapse for the points of interest along the SIDM halo evolution trajectory in Fig. 3. For the SIDM halo with $M_{200}=10^{11}~{}{\rm M_{\odot}}$, we find that our results align most closely with Q. Minor et al. (2021) at $\tau\simeq 1.84$. This indicates that an SIDM halo must evolve to be sufficiently core-collapsed to achieve results consistent with observations.
@@ 3.2 Mock lensing image test
In addition to comparing the compact, core-collapsed SIDM halo with observations, we also explore whether the overconcentrated halos observed in previous strong lensing studies could be explained by fitting a core-collapsed SIDM halo with a CDM halo profile.
To begin, we simulate a mock lensing image as a mimic of J0946+1006, as shown in Fig. 5. The macro model, which incorporates the mass and light distributions of the main lens and background source, along with their parameters, is based on the fitting results from Q. He et al. (2024). The lens mass is modeled as an elliptical power-law (EPL) distribution. As shown in Q. He et al. (2024), the lens light is well-represented by a multi-Gaussian expansion (MGE, M. Cappellari, 2002), so we exclude it from our mock image. For the background source, instead of using a parametric model, we adopt a pixelised flux distribution from Q. He et al. (2024).
The perturber is modeled as the SIDM halo that best matches the results of Q. Minor et al. (2021) (represented by the red star in Fig. 3). Directly calculating the deflection angle of a 3D density profile is challenging. To overcome this issue, we use the MGE model to fit the 3D density profile and then convert the fitting result into a surface density to calculate the perturbation. For a single Gaussian model, the conversion from a 3D profile to a 2D projected profile can be performed analytically with:
@@
where $\rho(r;\sigma)\propto\exp[-r^{2}/2\sigma^{2}]$ is the 3D Gaussian density profile and $\sigma$ is the standard deviation of the profile. As shown in the right panel of Fig. 5, the MGE model accurately reproduces the density profile of the input SIDM halo, with a relative error of $\lesssim 3\%$ within 50 arcsec. The input position of the perturber, marked by a red cross in the right panel, is determined from the power-law fitting results of J. W. Nightingale et al. (2024).
@@
Next, we fit the simulated mock lensing image. The lens mass is modeled using an EPL profile plus external shear. Following J. W. Nightingale et al. (2024), we model the source light using a Voronoi mesh with natural neighbor interpolation (R. Sibson, 1981). We smooth the source reconstruction by applying adaptive regularization based on the source fluxes at each point. The regularization formalism is the same as S. H. Suyu et al. (2006) and the likelihood function could be found in Eq. 17 of J. W. Nightingale et al. (2024). The positions of the Voronoi mesh centers on the image plane are adjusted to follow the morphology of the lensing arc, with more source pixels reconstructing the brighter regions. A detailed description of the adjustment scheme can be found in K. Wang et al. (2025).
For the subhalo, we adopt a tNFW profile:
@@
where $\rho_{s}$ and $r_{s}$ are the characteristic density and scale radius of the NFW profile (Eq. 4), and $r_{t}$ is the truncation radius. We fit the mock lensing image using the automated “SLaM” (Source, Lens, and Mass) strong lens modeling pipeline, which is implemented in the open-source software PyAutolens (X. Cao et al., 2022; A. Etherington et al., 2022; Q. He et al., 2023; J. W. Nightingale et al., 2024; Q. He et al., 2024). The model takes as input three parameters: the halo mass ($M_{200}$), concentration ($c_{200}$), and truncation radius ($r_{t}$), while $\rho_{s}$ and $r_{s}$ are subsequently derived from $M_{200}$ and $c_{200}$. We use the nested samplers Dynesty (J. S. Speagle, 2020) and Nautilus (J. U. Lange, 2023) to obtain the posteriors.
Figure 6: The posterior distributions of $M_{200}$, $c_{200}$ and $r_{t}$ from the tNFW model fitting. In the 1D posteriors, the black dashed lines indicate the medians and the $3\sigma$ boundaries. In the 2D posterior, the black contours represent the $1\sigma$, $2\sigma$, and $3\sigma$ regions, respectively. The red line marks the $M_{200}$ value of the input SIDM halo used to generate the mock image, while the orange dashed line corresponds to its $c_{200}$ value.
Figure 7: Same format as Fig. 2, but here we include the best-fit tNFW profile from our mock lensing image, with $M_{200}=2.40\times 10^{10}~{}{\rm M_{\odot}}$ and $c_{200}=62.25$. Left: 2D projected surface density profiles, with the gray band indicating the radius range used to calculate the log-slope, $\gamma_{\rm 2D}$. Right: 2D enclosed mass profiles. The error bar indicate the $3\sigma$ measurement uncertainty of $M_{\rm 2D,tot}~{}{\rm(<1~{}kpc)}$ from Q. Minor et al. (2021).
@@
In Fig. 6, we show the joint posterior distribution of $M_{200}$, $c_{200}$ and $r_{t}$ derived from the tNFW model fitting. The best-fit tNFW halo has a halo mass systematically lower than that of the input SIDM halo. Suppose that the $1\sigma$ scatter of concentration is 0.15 dex (J. Wang et al., 2020), the inferred concentration is extremely high, exceeding the CDM mass-concentration relation by $5.43\sigma$. This outcome is consistent with previous analyses of J0946+1006 (Q. Minor et al., 2021; D. J. Ballard et al., 2024; G. Despali et al., 2024; W. J. R. Enzi et al., 2024). Our results indicate that a model misspecification—incorrectly modeling a core-collapsed SIDM subhalo as a tNFW halo—leads to an overconcentrated solution.
Note that, since we do not know whether or not the dark perturber has undergone tidal stripping, we create the mock image using an untruncated SIDM halo as the input perturber. In practice, truncation has negligible impact on both the subhalo’s lensing effect (which primarily originates from the core region) and its inferred $M_{200}$ (which is essentially an extrapolation of the inner profile). As shown in Fig. 6, the observational data places no constraints on the tidal radius. Further discussion on the effects of tidal stripping can be found in Section 4.3.
For each sampled tNFW profile, we calculate its $\gamma_{\rm 2D}~{}{\rm(1~{}kpc)}$ and $M_{\rm 2D}~{}{\rm(<1~{}kpc)}$. We then compute the median and $1\sigma$ scatter of these 2D parameters and plot them in Fig. 3. One can find that the best-fit tNFW profile falls slightly out the $3\sigma$ region around the observed values. Fig. 7 further shows that, compared to the input SIDM halo, the inferred tNFW profile has a slightly flatter $\gamma_{\rm 2D}$ around 1 kpc and a lower $M_{\rm 2D}$ within 1 kpc. Although some differences exist, the tNFW profile is still able to capture the lensing effects produced by a core-collapsed halo. However, the fitting process tends to yield a lower halo mass than the true value while significantly overestimating the concentration parameter.

### v2

@@ 2.2 Computing the full time evolution of the SIDM halo
Figure 2: The density and mass profiles of a core-collapse SIDM halo with $M_{200}=10^{11}~{\rm M_{\odot}}$ (corresponding to the evolutionary stage indicated by the red star in Fig. 1) and its CDM counterpart. SIDM and CDM profiles are shown in blue and black, respectively. Left: 3D density profiles, with the blue vertical dashed-dotted line marking the stitching radius, $r_{1}$. Middle: 2D projected surface density profiles. Right: 2D enclosed mass profiles.
@@
In an SIDM halo, heat transport driven by collisions between dark matter particles governs the temporal evolution of halo internal structures. Initially, these collisions facilitate heat flow inward along a positive temperature gradient, leading to the formation of a central core. This process, known as core formation, results in the gradual growth of the core as the inner region undergoes thermalization. Once thermal contact extends to the outskirts of the halo, heat begins to flow outward, marking the onset of the core-collapse phase. Due to the negative specific heat characteristic, the core area becomes hotter and denser after losing heat, leading the temperature and density of the isothermal core at the halo center increase steadily, culminating in a gravothermal catastrophe (Tulin&Yu).
@@
The full time evolution of an SIDM halo can be more rigorously captured using the gravothermal fluid formalism, which models SIDM halos by solving a set of coupled partial differential equations for matter and heat transport with a CDM halo as the initial condition (e.g., Essig2019; Nishikawa2020; Yang2023). In contrast, the isothermal Jeans method assumes that the inner region is in approximate equilibrium, an assumption that breaks down during the core-collapse phase, making the standard approach inadequate for capturing the behavior of SIDM halos in this regime.
@@
However, Jiang2023 noted that prior to core collapse, the isothermal Jeans method allows for both high-density and low-density solutions for the isothermal core: during the minimization of $\delta^{2}$, two sets of solutions $(\rho_{0},\nu_{0})$ are found, where the $\nu_{0}$ values are close to each other, but the $\rho_{0}$ values differ by several orders of magnitude (see Figure 1 of Jiang2023). These two sets of solutions gradually converge at a specific time, denoted as $t_{\rm merge}$.
@@
Based on this, Yang2024 (Yang2024, inspired by private communications with M. Kaplinghat) demonstrated that mirroring the high-density solution at the merger time $t_{\rm merge}$ yields isothermal solutions corresponding to the core-collapse phase. This leads to a natural definition of the collapse timescale as $t_{\rm coll}=2\times t_{\rm merge}$. As shown in Figure 2 of Yang2024, although minor discrepancies exist, the gravothermal and isothermal solutions exhibit remarkable similarity throughout the entire halo evolution. The physical justification for this mirroring approach can be understood as follows. Since time is treated as a scalar quantity in Eq. 5, the formulation inherently lacks temporal directionality. This means the isothermal Jeans model invariably solves for density profiles that best approximate an isothermal sphere configuration at any given instant – even during core collapse. By mirroring the high-density solution about $t_{\rm merge}$, we essentially reconstruct the temporal sequence of this evolutionary phase while recovering the proper chronological ordering.
@@
In this work, we adopt the “mirror” strategy to calculate density profile of SIDM halo in core-collapse state. For a given CDM halo with known $M_{200}$ and $c_{200}$, we first compute the time $t_{\rm merge}\cdot\sigma_{m}$, in units of $\rm Gyr\cdot cm^{2}/g$}, when the high-density and low-density solutions merge. Then, by progressively increasing $t_{\rm age}\cdot\sigma_{m}$ values in Eq. 5 and solving for the density profile of the isothermal core, we determine which solution to adopt: if $t_{\rm age}\cdot\sigma_{m}<t_{\rm merge}\cdot\sigma_{m}$, the low-density solution is used; otherwise, the mirrored high-density solution is applied. This process is illustrated in Fig. 1 for a halo with $M_{200}=1\times 10^{11}~{\rm M_{\odot}}$, where the $c_{200}$ is calculated using the mass–concentration relation for CDM from Dutton2014.
@@
In the right panel of Fig. 1, we present the evolutionary scenario of the velocity dispersion, $\nu_{0}$, combining the low-density solution and the mirrored high-density solution, omitting the original high-density solution for clarity. After its initial increase, $\nu_{0}$ remains nearly constant for a long time before $t_{\rm merge}$ but increases steadily afterward. This behavior underscores the rationale for using $t_{\rm merge}$ as the starting point for the core-collapse phase in the isothermal Jeans method.
@@
In Fig. 2 we compare the density and mass profiles of a core-collapse SIDM halo ($M_{200}=10^{11}~{\rm M_{\odot}}$) with its CDM counterpart, highlighting differences in matter distribution. The corresponding evolutionary stage is marked by a red star in Fig. 1. The inner region of the SIDM halo develops a steep and compact core with a radius smaller than 1 kpc, while the density profiles of the SIDM and CDM halos converge at $r_{1}$. The core-collapse-induced inward mass aggregation results in the SIDM halo exhibiting higher densities than the CDM halo within the 0.03 to 3 kpc range. This steeper inner density profile enables the SIDM halo to accumulate more projected mass up to several kpc, with the increase in enclosed projected mass being most pronounced at 1 kiloparsecs. It also leads to a steeper slope of the surface density around 1 kpc, as illustrated in the middle and right panels of Fig. 2.
Readers can refer to Section 2 of Z. Jia et al. (2025, in preparation) for more detailed steps on how to solve $t_{\rm merge}$. Furthermore, we find that the semianalytical model introduced above provides a reliable approximation for the evolution of an SIDM halo. A detailed comparison (S. Li et al., in preparation) with a new high-resolution $N$-body simulation by Fischer2025 shows that our model accurately reproduces the density profile for over 95% of the simulation’s core-collapse evolution and is sufficiently accurate for the purposes of this study.
@@ 2.3 Computing the lensing properties
To compare with observations of J0946+1006, we calculate two lensing properties that, following Minor2021, have been identified as well constrained through lens modeling: (1) the average logarithmic slope of the surface density profile within 1 kpc, denoted as $\gamma_{\rm 2D}~(1~{\rm kpc})$, and (2) the projected total mass within 1 kpc, $M_{\rm 2D,tot}~(<1~{\rm kpc})$.
We adopt these two properties from the best-fit results of “tNFWmult” model in Minor2021, along with model results from Despali2024, as observational benchmarks for comparison with our SIDM model results. To assess whether the modeled SIDM halos produce observational signatures akin to those detected, and to determine the evolutionary stages at which these occur, we computed their complete time evolution and determined the density profiles at specific snapshots along their evolutionary paths, as illustrated by the black dots in Fig. 1.
@@
To measure $\gamma_{\rm 2D}$ at 1 kpc, we fit the surface mass density profile $\Sigma(R)$ with a power-law function $\Sigma(R)=\Sigma_{0}R^{\gamma}$ over the radial range 0.75–1.25 kpc, matching the analysis region of Minor2021. $M_{\rm 2D}~(<1~{\rm kpc})$ is then obtained by integrating $\Sigma(R)$ from 0.001 to 1 kpc along $R$.
@@ 3.1 Comparison with observations
In Fig. 3, we compare the lensing properties of SIDM halos, evaluated after $t_{\rm merge}$, with those of the observed dark perturber in J0946+1006 (Minor2021; Despali2024). We consider four halo masses in ascending order: $5\times 10^{9}$, $3\times 10^{10}$, $1\times 10^{11}$, and $5\times 10^{11}~M_{\odot}$. Their $c_{200}$ values are calculated using the mass–concentration relation from Dutton2014.
Our analysis indicates that, across all four selected halo masses, the matter condensation resulting from core collapse causes $\gamma_{\rm 2D}~(1~{\rm kpc})$ to reach approximately -1.4 at its steepest point, which is steeper than the singular isothermal sphere case, -1. This steepening brings the slope close to the value reported by Minor2021: $\gamma_{\rm 2D}~(1~{\rm kpc})=-1.27^{+0.11}_{-0.13}$. However, only the core-collapse halo with a mass of $10^{11}~{\rm M_{\odot}}$ achieves an enclosed mass comparable to the observed value of $(3.3\pm 0.3)\times 10^{9}~{\rm M_{\odot}}$. Halos with masses significantly larger or smaller than the optimal range accumulate either too much or too little mass in this phase, inconsistent with the observation constraints.
@@
Another important observation is that, for a given halo mass, $M_{\rm 2D}~{\rm(<1~kpc)}$ gradually converges to a finite value during the core-collapse stage, and after $\gamma_{\rm 2D}~{\rm(1~kpc)}$ reaches its turning point, it reverses accordingly. This is because the dark matter self-interaction is only effective at the inner region of the halo, and the redistributable mass during core collapse is constrained by the total mass of the dark halo within $r_{1}$. Therefore, for any given CDM halo mass, the properties of the density core created by SIDM are not arbitrary; they depend on the initial density profile of the halo at the beginning of its evolution.
@@
From another perspective, the degree of core collapse in the SIDM halo can be analyzed within the gravothermal fluid framework. In this context, radial heat transport, driven by dark matter self-interactions, can be characterized by the ratio of the scattering mean free path to the halo’s dynamical scale height. The time at which the innermost shell of an SIDM halo transitions to the short mean-free-path regime signifies the onset of post-self-similar collapse. At this point, the inner density begins to increase dramatically, making this time a defining collapse timescale, denoted as $t_{\rm c}$. Essig2019 performed numerical simulations and obtained numerical templates between the core-collapse timescale and other model parameters,
@@
where $\beta=0.75$ is calibrated with $N$-body simulations. Based on this scaling relation, we can transform the product $t_{\rm age}\cdot\sigma_{m}$ into a dimensionless time variable $\tau=t/t_{\rm c}$. This allows us to analyze the degree of core collapse for the points of interest along the SIDM halo evolution trajectory in Fig. 3. For the SIDM halo with $M_{200}=10^{11}~{\rm M_{\odot}}$, we find that our results align most closely with Minor2021 at $\tau\simeq 1.84$. This indicates that an SIDM halo must evolve to be sufficiently core-collapsed to achieve results consistent with observations.
@@ 3.2 Mock lensing image test
In addition to comparing the compact, core-collapse SIDM halo with observations, we also explore whether the overconcentrated halos observed in previous strong-lensing studies could be explained by fitting a core-collapse SIDM halo with a CDM halo profile.
To begin, we simulate a mock lensing image as a mimic of J0946+1006, as shown in Fig. 5. The macro model, which incorporates the mass and light distributions of the main lens and background source, along with their parameters, is based on the fitting results from He_2024. The lens mass is modeled as an elliptical power-law (EPL) distribution. As shown in He_2024, the lens light is well represented by a multi-Gaussian expansion (MGE; Cappellari_2002), so we exclude it from our mock image. For the background source, instead of using a parametric model, we adopt a pixelized flux distribution from He_2024.
The perturber is modeled as the SIDM halo that best matches the results of Minor2021. Directly calculating the deflection angle of a 3D density profile is challenging. To overcome this issue, we use the MGE model to fit the 3D density profile and then convert the fitting result into a surface density to calculate the perturbation. For a single Gaussian model, the conversion from a 3D profile to a 2D projected profile can be performed analytically with:
@@
where $\rho(r;\sigma)\propto\exp[-r^{2}/2\sigma^{2}]$ is the 3D Gaussian density profile and $\sigma$ is the standard deviation of the profile. As shown in the right panel of Fig. 5, the MGE model accurately reproduces the density profile of the input SIDM halo, with a relative error of $\lesssim 3\%$ within 50 arcsec. The input position of the perturber, marked by a red cross in the right panel, is determined from the power-law fitting results of Nightingale_24.
@@
Next, we fit the simulated mock lensing image. The lens mass is modeled using an EPL profile plus external shear. Following Nightingale_24, we model the source light using a Voronoi mesh with natural neighbor interpolation (Sibson1981). We smooth the source reconstruction by applying adaptive regularization based on the source fluxes at each point. The regularization formalism is the same as Suyu2006 and the likelihood function could be found in Eq. 17 of Nightingale_24. The positions of the Voronoi mesh centers on the image plane are adjusted to follow the morphology of the lensing arc, with more source pixels reconstructing the brighter regions. A detailed description of the adjustment scheme can be found in Wang_2025.
For the subhalo, we adopt a tNFW profile,
@@
where $\rho_{s}$ and $r_{s}$ are the characteristic density and scale radius of the NFW profile (Eq. 4), and $r_{t}$ is the truncation radius. We fit the mock lensing image using the automated Source, Lens, and Mass (SLaM) strong lens modeling pipeline, which is implemented in the open-source software PyAutolens (slam_Cao_2022; slam_Amy22; slam_he; He_2024; Nightingale_24). The model takes as input three parameters: the halo mass ($M_{200}$), concentration ($c_{200}$), and truncation radius ($r_{t}$), while $\rho_{s}$ and $r_{s}$ are subsequently derived from $M_{200}$ and $c_{200}$. We use the nested samplers Dynesty (dynesty) and Nautilus (nautilus) to obtain the posteriors.
Figure 6: The posterior distributions of $M_{200}$, $c_{200}$ and $r_{t}$ from the tNFW model fitting. In the 1D posteriors, the black dashed lines indicate the medians and the $3\sigma$ boundaries. In the 2D posterior, the black contours represent the $1\sigma$, $2\sigma$, and $3\sigma$ regions. The red line marks the $M_{200}$ value of the input SIDM halo used to generate the mock image, while the orange dashed line corresponds to its $c_{200}$ value.
Figure 7: Same format as Fig. 2, but here we include the best-fit tNFW profile from our mock lensing image, with $M_{200}=2.40\times 10^{10}~{\rm M_{\odot}}$ and $c_{200}=62.25$. Left: 2D projected surface density profiles, with the gray band indicating the radius range used to calculate the log slope, $\gamma_{\rm 2D}$. Right: 2D enclosed mass profiles. The error bar indicates the $3\sigma$ measurement uncertainty of $M_{\rm 2D,tot}~{\rm(<1~kpc)}$ from Minor2021.
@@
In Fig. 6, we show the joint posterior distribution of $M_{200}$, $c_{200}$ and $r_{t}$ derived from the tNFW model fitting. The best-fit tNFW halo has a halo mass systematically lower than that of the input SIDM halo. Supposing that the $1\sigma$ scatter of concentration is 0.15 dex (Wang2020), the inferred concentration is extremely high, exceeding the CDM mass–concentration relation by $5.43\sigma$. This outcome is consistent with previous analyses of J0946+1006 (Minor2021; Ballard_24; Enzi2024; Despali2024). Our results indicate that a model misspecification–incorrectly modeling a core-collapse SIDM subhalo as a tNFW halo–leads to an overconcentrated solution.
Note that, since we do not know whether or not the dark perturber has undergone tidal stripping, we create the mock image using an untruncated SIDM halo as the input perturber. In practice, truncation has negligible impact on both the subhalo’s lensing effect (which primarily originates from the core region) and its inferred $M_{200}$ (which is essentially an extrapolation of the inner profile). As shown in Fig. 6, the observational data places no constraints on the tidal radius. Further discussion of the effects of tidal stripping can be found in Section 4.3.
For each sampled tNFW profile, we calculate its $\gamma_{\rm 2D}~{\rm(1~kpc)}$ and $M_{\rm 2D}~{\rm(<1~kpc)}$. We then compute the median and $1\sigma$ scatter of these 2D parameters and plot them in Fig. 3. One can find that the best-fit tNFW profile falls slightly outside the $3\sigma$ region around the observed values. Fig. 7 further shows that, compared to the input SIDM halo, the inferred tNFW profile has a slightly flatter $\gamma_{\rm 2D}$ around 1 kpc and a lower $M_{\rm 2D}$ within 1 kpc. Although some differences exist, the tNFW profile is still able to capture the lensing effects produced by a core-collapse halo. However, the fitting process tends to yield a lower halo mass than the true value while significantly overestimating the concentration parameter.